OpenStudy (anonymous):

FAN AND MEDAL System A - -4x + 6y = -2 2x + y = 5 System B -10x + 19y = -1 -4x + 6y = -2

OpenStudy (btaylor):

Do you know how to do elimination?

OpenStudy (anonymous):

Which of the following statements is correct about the two systems of equations? They will have the same solutions because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 3 times the second equation of System A. The value of x for System B will be –5 times the value of x for System A because the coefficient of x in the first equation of System B is –5 times the coefficient of x in the first equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –12 to the first equation of System A and the second equations are identical.

OpenStudy (anonymous):

Sorry that took me so long, the connection was really slow /:

OpenStudy (btaylor):

I think in order to answer that you need to solve the systems.

OpenStudy (anonymous):

I don't really know how to do that.. I've had a hard time with this.

OpenStudy (btaylor):

In order to solve by elimination, you want to multiply one equation by some number so that one of the variables cancels out. Let's look at system A first and eliminate x. What should we multiply the second equation by so that, when you add the two equations, the x cancels itself out?

OpenStudy (anonymous):

I'm not sure, I'm sorry ugh. I'm good at algebra it's just for some reason this is extremely difficult for me. I don't want to waste your time!

OpenStudy (btaylor):

You're not wasting my time if you are willing to learn. What do you need to add to -4x to make 0?

OpenStudy (anonymous):

Thank you so much you have no idea. 4? But you can't add 4 to -4x because 4 doesn't have x also right?

OpenStudy (btaylor):

What about 4x?

OpenStudy (anonymous):

Yes. So then with the second equation would you subtract 2x ?

OpenStudy (btaylor):

No. What do you need to multiply 2x by to make 4x?

OpenStudy (anonymous):

2x ?

OpenStudy (btaylor):

no, just 2. So, if we multiply all of the second line by 2, we get 4x+2y=10 now, we add the two equations together:  So we get 8y=8. That makes y= ?

OpenStudy (anonymous):

1 ?

OpenStudy (anonymous):

because you divide by 8 on both sides to get y alone right?

OpenStudy (btaylor):

exactly. Now, we have to plug in that value of y into one of the equations in system A to find x.

OpenStudy (anonymous):

So if I did the first equation, Would it be -4x + 6 = -2 ? So solving that you would get x = 2 ?

OpenStudy (btaylor):

perfect. So your solution for system A is (2, 1)

OpenStudy (btaylor):

Now, for system B: What should we multiply the second equation by to eliminate x when we add it together?

OpenStudy (btaylor):

^ actually we will need to multiply both equations here.

OpenStudy (anonymous):

4x right?

OpenStudy (anonymous):

But for the first equation in system B wouldn't you add 10 ?

OpenStudy (btaylor):

We'll try to make -20x on top, so we multiply the top equation by ?? We'll try to make +20x on bottom, so we multiply the bottom equation by ??

OpenStudy (anonymous):

No wait that was wrong . Ugh I don't know. So we'll multiply the top equation by 2 ? And we'll multiply the bottom by 5 ?

OpenStudy (btaylor):

bottom by -5 (you have -4,so we want to cancel out the negative). But good job otherwise. So when we multiply each term by those, what are our new equations?

OpenStudy (anonymous):

19y = -1 6y = -2 ?

OpenStudy (btaylor):

Not quite. We are looking for: 2(-10x) + 2(19y) = 2(-1) AND -5(-4x) - 5(6y) = -5(-2). Can you simplify those?

OpenStudy (anonymous):

I'm not sure.. I don't think so because you cant simplify x and y right ?

OpenStudy (btaylor):

I meant multiply them out.

OpenStudy (anonymous):

I don't know. I'm so sorry.

OpenStudy (btaylor):

I'll do the first one... -20x + 38y = -2 Seeing what I did, can you do the next one?

OpenStudy (anonymous):

20x - 30y = 10?

OpenStudy (btaylor):

perfect! Now, doing the addition I did before, what do you get?

OpenStudy (anonymous):

if I'm adding the 30y and the 38y would the 30y be negative because it was being subtracted ? If so, the equation would be 8y = 8 So y = 1 again

OpenStudy (btaylor):

Yep! Now, can you solve for x?

OpenStudy (anonymous):

-10x + 19 = -1 -10x = -20 x = 2

OpenStudy (btaylor):

Yes, and that matches what we have before (2,1). Now going back to the 4 options, we have the same solution for both systems. So, which one is it?

OpenStudy (btaylor):

The options are phrased weirdly.

OpenStudy (anonymous):

A or B ?

OpenStudy (btaylor):

I don't understand why they would do it the way they are, but B is right.